phoebeR:

Which of the following is a polynomial with roots negative square root of 5, square root of 5, and 3? Can you explain how to solve this?

phoebeR:

the options are x3 + 2x2 − 3x − 6 x3 − 2x2 − 3x + 6 3 + 3x2 − 5x − 15 x3 − 3x2 − 5x + 15

supie:

\[y = (x + √5)(x - √5)(x - 3) \]\[y = (x² - 5)(x - 3) \]\[y = x³ - 3x² - 5x +15\] I think that's how you do it.

Ashoka:

Yeah that is how you do it, Nice explanation Supie! \[(x+\sqrt(5))(x-\sqrt(5)) = x^2-xsqrt(5)+xsqrt(5)-5\] and then he canceled the xsqrt(5) terms

ToxicaDennisIce:

yu

starbucksgirl13:

think